OpenStudy (anonymous):

Help me plox! \[(2+\frac{ 3 }{ x })^2-(2+\frac{3 }{ x })-12=0\]

OpenStudy (anonymous):

do you know quadratic formula?

OpenStudy (anonymous):

yea @study100

OpenStudy (anonymous):

\[x= \frac{ -b \pm \sqrt b^2-4ac}{ 2a }\]

OpenStudy (anonymous):

Yes, first, distribute everything out \[4+ \frac{ 12 }{ x} + \frac{ 9 }{ x ^{2} } + -2 - \frac{ 3 }{ x} -12 = 0\]

OpenStudy (anonymous):

Now you want to get rid of the fractions so multiply everything by x^2 \[x ^{2}(4+ \frac{ 12 }{ x} + \frac{ 9 }{ x ^{2} } + -2 - \frac{ 3 }{ x} -12 = 0)\]

OpenStudy (jdoe0001):

\(\bf \large \begin{array}{cccllll} a^2&-a&-12\\ &\uparrow &\uparrow \\ &-4+3&-4\cdot 3 \end{array}\)

OpenStudy (anonymous):

Which gives you : \[4^{2} +12 x + 9 - 2x ^{2} -3x -12x ^{2}\]

OpenStudy (anonymous):

Combine like terms

OpenStudy (anonymous):

= -10 x^2 + 9x + 9 \[= - (10 x ^{2} - 9x - 9)\] a = 10 b= -9 c=- 9

OpenStudy (anonymous):

sorry I meant \[4 x ^{2} + 12x + 9 - 2x^{2} -3x -12x^{2}\]

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

I overtyped that x under 4x^2

OpenStudy (anonymous):

yea i got you so its -10x^2+9x+9

OpenStudy (anonymous):

Yes, now you can use the quadratic formula, or take out negative sign - (10x^2 -9x -9) either way, same answer

OpenStudy (anonymous):

then i just plug into quadratic formula

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

x= 3/2, -3/5

OpenStudy (anonymous):

Yes :)